# Multi-Domain Evolutionary Optimization of Network Structures

### v1

@@ IV-A Graph similarity
The measurement of graph similarity involves the consideration of the interactions between nodes and higher-order structures, enabling us to perform selective knowledge transfer between networks. Instead of transferring all available information, it allows us to focus exclusively on pertinent domains, minimizing the risk of negative transfer and reducing the computational complexity.
In general, the graph similarity can be easily obtained byaveraging node similarity of networks [48] or computing the distance of graph embeddings [49, 50]. However, the former approach oversimplifies the complex interactions within the graph, ignoring the community structure and global patterns of networks. On the other hand, existing graph pooling techniques are typically trained on small graphs, potentially constraining their efficacy in accurately representing large-scale graphs.
In addition to overcoming the above problems, this similarity is also expected to facilitate the construction of training data for the network alignment model. To this end, we propose a community-based measurement of graph similarity. Given a network $G=(V,E)$ and its community distribution $\widetilde{\mathcal{C}}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\right\}$, for each community $\mathcal{C}\in\widetilde{\mathcal{C}}$, we can define $D_{\mathcal{C}}=\{d_{v}|v\in\mathcal{C}\}$ as the degree distribution of community $\mathcal{C}$ where $d_{v}$ is the node degree of $v$. With the consideration of the scales of different networks, we normalize this distribution by dividing the largest degree of the network, denoted as $D^{\prime}_{\mathcal{C}}=\{d_{v}/d^{max}|v\in\mathcal{C}\}$ where $d^{max}$ is the largest degree in the network $G$. In addition, another challenge emerges due to the varying lengths of these normalized distributions across networks, which complicates the direct comparison of community degree distributions. To address this issue, we introduce a new definition that unifies the degree distributions across communities of different scales into the same length.
@@
Degree interval-based distribution: Given a community $\mathcal{C}\in\widetilde{\mathcal{C}}$, then we can obtain the normalized degree distribution $D^{\prime}_{\mathcal{C}}=\left\{d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{n}\right\}$ where $n$ is the number of nodes in community. Let $I_{k}=\left[a_{k},b_{k}\right)$ represent the $k$th interval, where $a_{k}$ is the lower bound, and $b_{k}$ is the upper bound of the interval, then we can have
@@
×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|),$ (7)
where $\mathbb{KL}(.)$ refers to Kullback–Leibler divergence and the second item $\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|)$ is used to evaluate the difference of the scales of communities.
Traditional graph representations focus on pairwise relationships between nodes while high-order graphs extend this concept and it is able to capture interactions involving more than two nodes. Therefore, we also consider the high-order information of networks, allowing for a more refined comparison that captures structural patterns. Given a network $G$ and a motif $M$, the network $G$ can be converted into a weighted network $G_{M}$ as follows:
@@
where ${E}^{M}\in E$ represents the set of edges within ${G}^{M}$. We can derive the adjacency matrix of hypergraph, denoted as $A_{M}$, in which each entry $(A_{M})_{ij}$ corresponds to the count of motifs that include the edge $e_{ij}$. We can measure the high-order difference between $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as follows:
@@
×$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|).$ (9)
The similarity between communities is then defined as follows:
$\mathcal{S}(\mathcal{C},\mathcal{C}^{\prime})=\textbf{Exp}(-\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right))+\textbf{Exp}(-\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right))$ (10)
@@
Aligned community: Given two networks $G_{A}$, $G_{B}$ and their corresponding community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, we can have the community-wise similarity matrix $\mathcal{S}\in\mathbb{R}^{k\times k^{\prime}}$ where $\mathcal{S}_{i,j}=\mathcal{S}(\mathcal{C}_{i},\mathcal{C}^{\prime}_{j}),i\in\{1,2,\cdots,k\},j\in\{1,2,\cdots,k^{\prime}\}$ Then, we define the set of aligned communities as follows:
$\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\{\mathcal{C}_{a_{1}},\mathcal{C}^{\prime}_{b_{1}}\},\{\mathcal{C}_{a_{2}},\mathcal{C}^{\prime}_{b_{2}}\},\cdots,\{\mathcal{C}_{a_{z}},\mathcal{C}^{\prime}_{b_{z}}\}\},$ (11)
where $z=\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)$. $\mathcal{C}_{a_{i}}$ and $\mathcal{C}^{\prime}_{b_{i}}$ refer to the $a_{i}$th and $b_{i}$th community in $\widetilde{\mathcal{C}_{A}}$ and $\widetilde{\mathcal{C}_{B}}$, respectively. Accordingly, we can have the set of their corresponding similarity values as follows:
$\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}=\{\mathcal{S}_{a_{1},b_{1}},\mathcal{S}_{a_{2},b_{2}},\cdots,\mathcal{S}_{a_{z},b_{z}}\},$ (12)
@@
$\mathcal{S}^{\mathcal{G}}_{A,B}=\frac{1}{\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)}\sum_{x\in\mathcal{S}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}}x.$ (14)
The process of calculating the graph similarity can be found in Algorithm 1.
@@
Output: The quantitative similarity $S_{A,B}$ of $G_{A}$ and $G_{B}$.
1:Obtain high-order graphs ${G}^{m}_{A}=({V}_{A},{E}^{m}_{A})$ and ${G}^{m}_{B}=(V_{B},{E}^{{m}}_{B})$
@@
7:     Find the largest value of $\mathcal{S}$ and its two corresponding communities as well as its index $i_{\text{min}},j_{\text{min}}=\text{argmin}S_{i,j}$ where $1\leq i\leq k,1\leq j\leq k^{\prime}$
@@
9:     Make the $i_{\text{min}}$ row and $j_{\text{min}}$ column of $S$ as 00
@@
11:Normalize the similarity $\mathcal{S}^{\mathcal{G}}_{A,B}$ based on the number of communities of $G_{A}$ and $G_{B}$
@@ IV-B Graph embedding
To construct the node mappings across different networks, we employ graph autoencoders (GAE) to generate embeddings since it can capture meaningful and compact graph representations without the labeled data [51]. In GAE, the graph convolutional network (GCN) is taken as the encoder [52], i.e.,
@@
where $\mathbf{A}$ denotes the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ denotes nodes features. The information aggregation of GCN is defined as follows:
@@
The decoder takes the latent space representation $\mathbf{Z}$ as the input to reconstruct the adjacency matrix. The inner-product is adopted as the decoder function:
@@ IV-C Network alignment
In the supervised learning component, a set of anchor nodes is carefully selected in terms of the structural role in both networks, serving as labels to train the network alignment model. Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, we can then obtain their aligned community pairs $\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$ as in Definition 2, which is the basis of the construction of anchor nodes. To make the network alignment model adapted to the cross-domain scenario, we consider both the important nodes and trivial nodes in each aligned community pair as anchor nodes.
@@
Anchor node (large degree) – $\mathcal{A}^{large}$: Let $v^{X,l}_{i}$ denote the $i$th largest-degree node in node set $X$. Then, given two networks $G_{A}$ and $G_{B}$, and an aligned community pair $\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}$, the set of anchor node of large degree is denoted as follows:
$\mathcal{A}^{large}=\bigcup\limits^{\{\mathcal{C},\mathcal{C}^{\prime}\}\in\widetilde{\mathcal{C}_{A}},\widetilde{\mathcal{C}_{B}}}\{v^{\mathcal{C},l}_{i},v^{\mathcal{C}^{\prime},l}_{i}\},i\in\{1,...,k_{l}\},$ (19)
@@
$\begin{split}\mathcal{L}_{large}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{large}}dist(\Phi^{a\rightarrow b}(\mathbb{E}(u)),\mathbb{E}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}(v)),\mathbb{E}(u)).\end{split}$ (20)
where $dist$ refers to mean square error (MSE), measuring the difference between two embeddings.
@@
Anchor node (small degree) – $\mathcal{A}^{small}$: Let $v^{X,s}_{i}$ denote the $i$th least-degree node in the node set $X$. Given an aligned pair of nodes $\{u,v\}\in\mathcal{A}^{large}$, we can define
@@
where $k_{s}=\mathbf{Min}\{log_{2}|\mathcal{N}(u)|,log_{2}|\mathcal{N}(v)|\}$ and $\mathcal{N}(.)$ refers to the neighbor nodes.
@@
$\mathcal{L}_{alignment}\left(\mathbf{W}_{s,t},\mathbf{b}_{s,t},\mathbf{W}_{t,s},\mathbf{b}_{t,s}\right)=\mathcal{L}_{large}+\mathcal{L}_{small}+\mathcal{L}_{us}.$ (24)
Upon the optimization on $\mathcal{L}_{alignment}$, the trained mapping can align the remaining nodes between the two networks, enabling the solutions between networks to be transferred to each other.
@@
Output: Graph Mapping model $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$
@@
4:     Map the embedding $\mathbb{E}_{A}(u)$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$, i.e., $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then map the embeddings of $G_{B}$ back to the space of $G_{A}$ via $\mathbb{M}_{BA}$
@@
6:     Map the embedding $\mathbb{E}_{B}(v)$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$, i.e., $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and then map the embeddings of $G_{A}$ back to the space of $G_{B}$ via $\mathbb{M}_{AB}$
@@
8:     Calculate the unsupervised loss between the original embeddings and the automapped embeddings.
@@
10:     Map $u$ to the space of $G_{B}$ via $\mathbb{M}_{AB}$ to obtain $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then calculating its difference with $\mathbb{E}_{A}(u)$
11:     Map $v$ to the space of $G_{A}$ via $\mathbb{M}_{BA}$ to obtain $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$ and then calculating its difference with $\mathbb{E}_{B}(v)$
@@
14:Update the parameters of mappings $\mathbb{M}_{AB}$ and $\mathbb{M}_{BA}$

### v2

@@ IV-A Graph similarity
The measurement of graph similarity enables us to perform selective knowledge transfer between networks. Instead of transferring all available knowledge, it allows us to focus exclusively on pertinent domains, minimizing the risk of negative transfer and reducing computational complexity.
In general, the graph similarity can be easily obtained by averaging node similarity of networks [57] or computing the distance of graph embeddings [58, 59]. However, the former approach oversimplifies the complex interactions within the graph, ignoring the community structure and global patterns of networks. On the other hand, existing graph pooling techniques are typically trained on small graphs, potentially constraining their efficacy on large-scale graphs.
Here, we propose a community-level measurement of graph similarity. Given a network $G=(V,E)$ and its community distribution $\widetilde{\mathcal{C}}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\right\}$, for each community $\mathcal{C}\in\widetilde{\mathcal{C}}$, we can define $D_{\mathcal{C}}=\{d_{v}|v\in\mathcal{C}\}$ as the degree distribution of community $\mathcal{C}$ where $d_{v}$ is the node degree of $v$. With the consideration of the scales of different networks, we normalize this distribution by dividing the largest degree of the network, denoted as $D^{\prime}_{\mathcal{C}}=\{d_{v}/d^{max}|v\in\mathcal{C}\}$ where $d^{max}$ is the largest degree in the network $G$. In addition, another challenge emerges due to the varying lengths of these normalized distributions across networks, which complicates the direct comparison of community degree distributions. To address this issue, we unify the degree distributions across communities of different scales into the same length.
@@
Degree interval-based distribution: Given a community $\mathcal{C}\in\widetilde{\mathcal{C}}$, then we can obtain the normalized degree distribution $D^{\prime}_{\mathcal{C}}=\left\{d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{n}\right\}$ where $n$ is the number of nodes in community. Let $I_{k}=\left[a_{k},b_{k}\right)$ represent the $k$-th interval, where $b_{k}$ is the upper bound of the interval and $a_{k}$ is the lower bound, then we can have
@@
$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|),$ (7)
where $\mathbb{KL}(.)$ refers to the Kullback–Leibler divergence and the second item $\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|)$ is used to quantify the disparity in community sizes.
Traditional graph representations focus on pairwise relationships between nodes while higher-order graphs extend this concept and can capture interactions involving more than two nodes. Therefore, we also take into account higher-order network knowledge, enabling a more refined analysis of structural information. Given a network $G$ and a motif $M$, the network $G$ can be converted into a weighted network $G_{M}$ as follows:
@@
where ${E}^{M}\in E$ represents the set of edges of ${G}^{M}$. We can measure the higher-order difference between $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as follows:
@@
$\quad\times\mathbf{Max}(|\mathcal{C}|/|\mathcal{C}^{\prime}|,|\mathcal{C}^{\prime}|/|\mathcal{C}|).$ (9)
The similarity between communities is defined as follows:
$\mathcal{S}(\mathcal{C},\mathcal{C}^{\prime})=\textbf{Exp}(-\operatorname{Diff}\left(\mathcal{C},\mathcal{C}^{\prime}\right))+\textbf{Exp}(-\operatorname{Diff}^{M}\left(\mathcal{C},\mathcal{C}^{\prime}\right)).$ (10)
@@
Aligned community: Given two networks $G_{A}$, $G_{B}$ and their corresponding community structures $\widetilde{\mathcal{C}_{A}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}_{B}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, we can have the community-wise similarity matrix $\mathcal{S}_{i,j}\in\mathbb{R}^{k\times k^{\prime}}$ where $i\in\{1,2,\cdots,k\},j\in\{1,2,\cdots,k^{\prime}\}$. Then, we define the set of aligned communities as follows:
$\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}=\{\{\mathcal{C}_{a_{1}},\mathcal{C}^{\prime}_{b_{1}}\},\{\mathcal{C}_{a_{2}},\mathcal{C}^{\prime}_{b_{2}}\},\cdots,\{\mathcal{C}_{a_{z}},\mathcal{C}^{\prime}_{b_{z}}\}\},$ (11)
[…]
$\mathcal{S}^{\widetilde{\mathcal{C}}}_{A,B}=\{\mathcal{S}_{a_{1},b_{1}},\mathcal{S}_{a_{2},b_{2}},\cdots,\mathcal{S}_{a_{z},b_{z}}\},$ (12)
@@
$\mathcal{S}^{\mathcal{G}}_{A,B}=\frac{1}{\mathbf{Min}(|\widetilde{\mathcal{C}_{A}}|,|\widetilde{\mathcal{C}_{B}}|)}\sum_{x\in\mathcal{S}^{\widetilde{\mathcal{C}}}_{A,B}}x.$ (14)
The graph similarity matrix $\mathcal{S}^{\mathcal{G}}$ is initialized based on Eq. 14. Initially, it is used to identify the assisted network for each network and is subsequently updated to manage the transfer of solutions from each assisted network. The process of calculating the graph similarity is presented in Algorithm 1.
@@
Output: The graph similarity between $G_{A}$ and $G_{B}$
1:Obtain higher-order graphs ${G}^{M}_{A}=({V}_{A},{E}^{M}_{A})$ and ${G}^{M}_{B}=(V_{B},{E}^{{M}}_{B})$
@@
7:  Find the largest value of $\mathcal{S}$ and its two corresponding communities as well as its index $i_{\text{max}},j_{\text{max}}=\text{argmax}S_{i,j}$ where $1\leq i\leq k,1\leq j\leq k^{\prime}$
@@
9:  Set the $i_{\text{max}}$ row and $j_{\text{max}}$ column of $S$ to 0
@@
11:Normalize the similarity $\mathcal{S}^{\mathcal{G}}_{A,B}$
@@ IV-B Graph embedding
To construct the node mappings across different networks, we employ graph autoencoders (GAE) to generate embeddings capturing meaningful and compact graph representations without the labeled data [60]. In GAE, the graph convolutional network (GCN) is taken as the encoder [61], i.e.,
@@
where $\mathbf{A}$ denotes the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ denotes nodes features. The information aggregation of GCN is
@@
The decoder takes the latent space representation $\mathbf{Z}$ as input to reconstruct the adjacency matrix. The inner product is adopted as the decoder function:
@@ IV-C Network alignment
A collection of anchor nodes is carefully chosen based on their structural functions in both networks, serving as labels to train the network alignment model. Given two networks $G_{A}=(V_{A},E_{A})$ and $G_{B}=(V_{B},E_{B})$, we can then obtain their aligned community pairs $\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$ as in Definition 2, which is the basis of the construction of anchor nodes. To adapt the network alignment model for multi-domain scenarios, both trivial and important nodes within each aligned community pair will be considered as anchor nodes.
@@
Anchor node (large degree) – $\mathcal{A}^{large}$: Suppose the $i$-th largest-degree node in the set $X$ is $v^{X,l}_{i}$. Then, given two networks $G_{A}$ and $G_{B}$, and an aligned community pair $\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$, the set of anchor large-degree nodes can be formulated as follows:
$\mathcal{A}^{large}=\bigcup\limits^{\{\mathcal{C},\mathcal{C}^{\prime}\}\in\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}}\{v^{\mathcal{C},l}_{i},v^{\mathcal{C}^{\prime},l}_{i}\},i\in\{1,...,k_{l}\},$ (19)
@@
$\begin{split}\mathcal{L}_{large}&=\sum\limits_{\{u,v\}\in\mathcal{A}^{large}}dist(\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u)),\mathbb{E}_{B}(v))\\ &\quad+dist(\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v)),\mathbb{E}_{A}(u)),\end{split}$ (20)
where distdist refers to the mean square error (MSE), measuring the difference between two embeddings.
@@
Anchor node (small degree) – $\mathcal{A}^{small}$: Suppose the $i$-th least-degree node in the set $X$ is $v^{X,s}_{i}$. Given a set of aligned large-degree nodes $\{u,v\}\in\mathcal{A}^{large}$, the aligned small-degree nodes can be defined as follows:
@@
where $k_{s}=\mathbf{Min}\{log_{2}|\mathcal{N}(u)|,log_{2}|\mathcal{N}(v)|\}$ and $\mathcal{N}(.)$ denotes the neighbor nodes.
@@
$\mathcal{L}_{alignment}\left(\mathbf{W}_{a,b},\mathbf{b}_{a,b},\mathbf{W}_{b,a},\mathbf{b}_{b,a}\right)=\mathcal{L}_{large}+\mathcal{L}_{small}+\mathcal{L}_{us}.$ (24)
After optimizing $\mathcal{L}_{alignment}$, the trained mapping will align the remaining nodes between the two networks, facilitating the transfer of solutions between them.
To establish node-level alignment between two networks based on their learned embeddings, we first project the embeddings from one network into the latent space of the other using learned mapping functions. Specifically, for each node $u\in V_{A}$, its embedding is transformed into the space of $G_{B}$ using the mapping function $\Phi^{A\rightarrow B}$. The node $v\in V_{B}$ whose embedding is closest, in terms of Euclidean distance, to the mapped embedding of node $u$ is identified, and its index is stored in the alignment mapping as $M_{A\rightarrow B}(u)=v$.
@@
Output: Graph Mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$
@@
4:  Map the embedding $\mathbb{E}_{A}(u)$ to the space of $G_{B}$ via $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and then map the embeddings of $G_{B}$ back to the space of $G_{A}$
@@
6:  Map the embedding $\mathbb{E}_{B}(v)$ to the space of $G_{A}$ via $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and then map the embeddings of $G_{A}$ back to the space of $G_{B}$
@@
8:  Calculate the unsupervised loss between the original embeddings and the automapped embeddings
@@
10:  Map $u$ to the space of $G_{B}$ to obtain $\Phi^{a\rightarrow b}(\mathbb{E}_{A}(u))$, and calculating its difference with $\mathbb{E}_{B}(v)$
11:  Map $v$ to the space of $G_{A}$ to obtain $\Phi^{b\rightarrow a}(\mathbb{E}_{B}(v))$, and calculating its difference with $\mathbb{E}_{A}(u)$
@@
14:Update the parameters of mappings $\Phi^{a\rightarrow b}$ and $\Phi^{b\rightarrow a}$